OpenStudy (anonymous):

painter was hired to design and draw a mural for a new restaurant. before he can design the mural, he needs to know the are of the wall. the restaurant manager informs the painter that the mural wall is 6-√2 feet tall by 12-√2 feet wide. what is the total area of the mural wall? (recall that the area of a rectangle is the width multiplied by the length) a. 70-18√2 square feet b. 74-18√2 square feet c. 74-6√2 square feet d. 74+18√2 square feet

OpenStudy (anonymous):

is this \[6\sqrt{2}\] or \[6-\sqrt{2}\]??

OpenStudy (anonymous):

the second one .

OpenStudy (anonymous):

then your job is \[(6-\sqrt{2})(12-\sqrt{2})\] who measures a wall this way??

OpenStudy (anonymous):

btw answer is "b"